The most recent financial statements for Live Co. are shown here: Balance Sheet $4,677 Debt 11,450 Equity $16,127 Total IncomMultiple Choice 2.38% 3.85% 9.93% 4,05% 3.95%
Answer #1

ROA=Net income/Total assets

=(987/16127)

=0.061201711

Retention ratio=1-payout ratio

=(1-038)=0.62

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.061201711*0.62)/[1-(0.061201711*0.62)]

which is equal to

=3.95%(Approx).